OpenStudy (anonymous):

Identify the vertex, focus, and directrix. y=(1/12)(x^2)

OpenStudy (anonymous):

Ok so the normal equation of a parabola is \[y^2=4ax\] Which would be for a horizontal parabola with the focus on the x axis However the x^2 tells us that the parabola is U shaped so the focus is on the y axis Rearrange to the form x^2 = 4ay x^2 = 12y From this we can see that a= 3 Focus of this parabola is on the y axis so is at (0,a) = (0,3) Directrix y= -a y=-3 Parabola is centred on (0,0) So vertex is just that.

OpenStudy (anonymous):